Not logged in
Mon Apr 20 17:49:58 2020
Empty N2 replaced.
Logged back into zita and started the StripTool traces (even though we keep the TV off nowadays).
c1susaux acro-crate power cycled to re-enable PRM suspension control (all other vertex optics also now respond to slow bias voltage sliders being moved).
c1iscaux needed a hard reboot as it wasn’t seen on martian. I power cycled the crate for good measure.
Marconi turned back on with correct frequency/amplitude.
c0rga is now seen again on martian network. I re-enabled the RGA scanning so that it takes a scan every morning at 4am.
The forepumps for TP2/TP3 are noisier than I remember. The former has ~10,000 hrs on the clock. How often does the tip seal replacement need to happen?
HV supplies for ASX/ASY PZTs re-energized.
IFO re-aligned for locking.
c1oaf and c1daf models restarted. c1oaf required the usual start/stop/start sequence to make the DAQ errors go away, and luckily the FE didn’t crash when the model was unloaded.
POX/POY/PRMI 1f carrier/green locking all was smooth.
For some reason, the PRC angular FF filters i trained no longer do anything good
(but MCL is still good). collected 20mins of PRMI 1f locked data for investigations.
Update 21 Apr 2020 1200:
Looking at Attachments #1 and #2, the spectra for motion sensed by the POP QPD does indeed look very different on Apr 6 vs Apr 20. Could be some interference from Oplev loop or maybe some EPICS values didn't get reset correctly, needs more investigation. It doesn't seem reasonable to me that the plant changes by so much (spectra were taken at similar times of the day, ~5pm).
Update 22 Apr 2020 1500:
As suspected, the PRM oplev was disabled for whatever reason. Re-enabling it, I recovered the good performance from two weeks ago. ✅
Uploaded Tue Apr 21 12:47:22 2020
Uploaded Tue Apr 21 12:47:34 2020